# Towards Human-AI Synergy in UI Design: Enhancing Multi-Agent Based UI Generation with Intent Clarification and Alignment

### v1

@@ Abstract.
In automated user interface (UI) design generation, a key challenge is the lack of support for iterative processes, as most systems only focus on end-to-end generation of designs as starting points. This results from (1) limited capabilities to fully interpret user design intent from text or images, and (2) a lack of transparency, which prevents designers from refining intermediate results. To address existing limitations, we introduce PrototypeAgent, a human-centered, multi-agent system for automated UI generation. The core of PrototypeAgent is a theme design agent that clarifies implicit design intent through prompt augmentation, coordinating with specialized sub-agents to generate specific components. Designers interact with the system via an intuitive interface, providing natural language descriptions and layout preferences. During generation, PrototypeAgent enables designers to refine generated intermediate guidance or specific components, ensuring alignment with their intent throughout the generation workflow. Evaluations through experiments and user studies show PrototypeAgent’s effectiveness in producing high-fidelity prototypes that accurately reflect design intent as well as its superiority over baseline models in terms of both quality and diversity.
Interactive design, User interface, Large language models, Image generation, Design Intent Clarification, Multi-Agent, Intent-Design Alignment, Automated Prototype Generation
copyright: nonecopyright: acmcopyrightjournalyear: 2022doi: XXXXXXX.XXXXXXXbooktitle: The 37th Annual ACM Symposium on User Interface Software and Technology(UIST’ 24)price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and tools
Figure 1. An overview of the interactive user interface (UI) prototype creation process facilitated by PrototypeAgent, our multi-agent generation system. (A) Users input their design intentions through prompts and initial UI layouts. (B) Using our advanced multi-modal, multi-agent system, we produce detailed prototypes. (C) These prototypes offer a top-down perspective, allowing users to interactively refine content, ranging from themes to specific components. (D) The output is a high-fidelity (both visual and functional) prototype.
@@ 1. Introduction
To gain deeper insights into the current human-AI collaboration challenges in the UI design domain, we interviewed 10 professional UI/UX designers. This revealed five key shortcomings in the current interaction between humans and AI-driven design tools, including the lack of: (F1) seamless and flexible integration of knowledge bases, (F2) Design modality requirements for accurately expressing functional design intent with AI, (F3) Implicit design intent clarification to enhance design descriptions, (F4) Decoupling the generation process for human-centered intent-design alignment, and finally (F5) Ensuring design coherence throughout the process.
@@
In response to these observations, we introduce PrototypeAgent, a novel interactive design system that enables multi-agent collaboration for UI prototype generation with intent clarification and alignment. This system empowers designers to craft high-fidelity prototypes that are editable, customisable and comprehensible at each stage of the generation process. As depicted in Figure 1, PrototypeAgent begins with two design inputs: a textual description and a wireframe layout (Figure 1 (A)). The text outlines the general design requirements (e.g., “Starting page for PrototypeAgent: an intelligent design assistant”), while the wireframe offers a preliminary UI layout, enabling designers to convey their initial and nuanced design concepts.
@@
In the prototype generation phase (Figure 1(B)), PrototypeAgent employs a top-down strategy, beginning with a central agent that establishes the overarching theme and performs implicit intent clarification through prompt augmentation. Following the decoupling approach, the central agent coordinates the automatic, iterative creation of components, leveraging a cache pool to seamlessly integrate details from various sub-agents, each specially designed with expertise in different aspects. These specialized agents, including a text content generation agent, image content generation agent and retrieval-based icon agent, ensure alignment with the overall design. This integrated approach allows PrototypeAgent to maintain aesthetic consistency, elevating the quality and coherence of the final prototype.
@@
In the review and editing phase (Figure 1 (C)), PrototypeAgent offers deeper insights and detailed control within its interactive environment. It facilitates precise, iterative refinement, balancing automated generation with human customisation for efficient and personalized integration. Additionally, the generated prototypes (Figure 1 (D)) can be saved in SVG or JSON formats, highlighting the system’s practicality and supporting high-quality creative output.
@@
For evaluation, we first conducted two automated quantitative evaluations, which validate PrototypeAgent’s ability to generate realistic and detailed UI designs. Our ablation study further highlights the significant impact and necessity of each module. In addition to the automated evaluations, we carried out two user studies involving 16 participants to assess the perceived usefulness of our work. The findings suggest that PrototypeAgent was positively received and showed significant potential in addressing the five identified challenges. Our work includes three main contributions:
@@
  • To the best of our knowledge, this work is among the first to report findings from semi-structured interviews with professional UI/UX designers working with AI-powered design tools, detailing their current workflows and identifying five key shortcomings.
  • We introduce PrototypeAgent, a novel multi-agent prototype generation system that addresses implicit clarification through design prompt augmentation and ensures intent-design alignment and design coherence by decoupling the generation process. It transforms user descriptions and low-fidelity wireframes into detailed prototypes based on our specialized domain-specific knowledge bases, with user feedback integrated at each stage to ensure precision.
  • We conducted both quantitative and qualitative evaluations to confirmed the effectiveness of our human-in-the-loop prototype generation approach and demonstrated its potential to enhance the UI/UX design process.
@@ 2. Related Work
Adobe Illustrator Figma / Adobe XD Midjourney / Stable Diffusion Vercel’s V0 Uizard PrototypeAgent (Our)
@@
Implicit Intent Clarification N/A N/A Not Supported Not Supported Not Supported Supported (through the Theme Design Agent prompt augmentation)
Human-Centered Intent-Design Alignment N/A N/A Not Supported Not Supported Not Supported Supported (through generation decoupling and explicit intermediate guidance)
Design Coherence Among Elements (Semantic + Style) Manual Manual Implicit Unknown Unknown Explicit (through decoupling and multi-agent explicit generation)
@@ 2.1. GUI Design Tools and Techniques
For creative design inspiration, text-based GUI retrieval, such as Guigle(Bernal-Cárdenas et al., 2019), leverages automated crawling and natural language processing to perform efficient searches through app hierarchies. Systems like Gallery D.C.(Chen et al., 2019), GUI2WiRe(Kolthoff et al., 2020), and RaWi(Kolthoff et al., 2023a) have further enhanced component extraction from screenshots, enabling flexible search based on dimensions, color, and text. Visual and multi-modal retrieval methods extend this by incorporating richer inputs such as screenshots and wireframes. Methods like WAE(Chen et al., 2020a), VINS(Bunian et al., 2021), and Swire(Huang et al., 2019) bridge the gap between low and high-fidelity designs, while Screen2Vec(Li et al., 2021) introduces multimodal embeddings for diverse GUI content. In high-fidelity prototyping, tools like GUIGAN(Zhao et al., 2021) and research by Forrest et al.(Huang et al., 2021) focus on component retrieval and arrangement for detailed prototypes. However, these retrieval-based methods still rely heavily on existing designs, limiting creative flexibility. They address early-stage inspiration but leave much of the manual effort required for design refinement and customization.
@@
On the generative front, techniques like Layout2Image(Biswas et al., 2021), VAE(Bengio and LeCun, 2014), MidJourney(Midjourney, 2024), and Stable Diffusion(Rombach et al., 2022b) offer new possibilities for design generation. However, these methods often result in unstructured, non-editable outputs, making them challenging for GUI design. Projects such as PLay(Cheng et al., 2023) and DocSynth(Biswas et al., 2021) have made contributions to layout generation, facilitating the creation of low-fidelity prototypes. However, they provide limited support for detailed component creation or high-fidelity refinement. Industry tools like Uizard(Uizard, 2024), Vercel’s V0(Vercel, 2024), and Figma plugins(uxp, 2024) provide automated prototype generation, yet they primarily offer starting points. The missing puzzle piece is how AI can continually collaborate with designers throughout the design iteration workflow, from concept to refinement.
Our work addresses these limitations by making Implicit Intent Clarification throughout the generation process. Designer prompts are often vague, and the AI generation process is opaque, making it hard for designers to refine outputs effectively. Clarification is needed in two directions: (1) helping designers provide more precise inputs and (2) making the AI’s decision-making process transparent to allow better control and modification. As reviewed in Section 3 and Table 1, our approach bridges this gap by ensuring intent clarification before the entire generation and decoupling content generation into a more controlled manner.
2.2. Interactive Applications with Large Language Models
Large Language Models (LLMs) have significantly influenced HCI, showcasing remarkable capabilities in zero/few-shot learning and conversational interactions. These models have been instrumental in various applications, enhancing user experience and efficiency. Wang et al.(Wang et al., 2023a) employ LLMs to facilitate versatile conversational interactions within mobile UIs, a notable step in making user interfaces more accessible and intuitive. Stylette(Kim et al., 2022) innovatively integrates LLMs to interpret natural language commands for web design modifications, effectively translating them into corresponding CSS properties.
Further extending the LLMs’ utility, MenuCraft (Kargaran et al., 2023) demonstrates their effectiveness in generating menu designs, while SUGILITE (Li et al., 2017) leverages them for automating tasks on smartphones. Duan et al.(Duan et al., 2024) have explored LLMs’ potential in providing automatic feedback for UI designs based on guidelines, indicating the model’s analytical strength. Similarly, CollabCoder(Gao et al., 2024b) utilizes LLMs for collaborative qualitative analysis, emphasizing their role in team-based settings. Moreover, VIVID (Choi et al., 2024) presents an application of LLMs in generating vicarious dialogues from lecture videos, showcasing their adaptability in educational contexts.
@@ 2.2. Interactive Applications with Large Language Models
Some previous works have explored agents that clarify intents during the process of human-AI collaboration through interactions between end users and AI using natural language instructions (Li et al., 2018; Srivastava et al., 2017). SOVITE (Li et al., 2020) expanded on this by enhancing system transparency through a mutual disambiguation pattern (Oviatt, 1999), where inputs from one modality help clarify inputs from another, allowing for breakdown repair. However, the interactive design generation process remains under-explored, as it is often unreliable and confusing. Motivated by SOVITE and related insights, and as shown in Table 1, our decoupling generation method with Human-Centered Intent-Design Alignment addresses this gap by allowing both the designer and AI to collaboratively understand and edit intermediate guidance and results, improving clarity and disambiguation throughout the process.
2.3. Large Language Models for Agent
The emergent capabilities of LLMs, such as zero-shot prompting and complex reasoning (Wu et al., 2022a; Vaswani, 2017; Brown, 2020), have spurred their integration into advanced agent research (Betz et al., 2021; Liu et al., 2021; Lu et al., 2021; Wu et al., 2022b). Notable examples include ReAct’s (Yao et al., 2022) use of chain-of-thought prompting (Feng and Chen, 2024) for generating reasoning traces and task-specific actions, and tools like AutoGPT (Gravitas, 2023), Chameleon (Lu et al., 2024), and VOYAGER (Wang et al., 2023b) make use of their internal knowledge bases to automate tasks by formulating a curriculum of multiple subgoals to achieve a high-level goal.
These LLMs have shown exceptional conversational engagement, decision-making, and task-completion abilities, significantly improving communication between users and agents. In the HCI research, applications like AudiLens (Park and Choi, 2023) have developed intelligent agents for public speech practice simulations, and GestureGPT (Zeng et al., 2023) has pioneered a dual-agent gesture understanding framework. Dang et al.(Dang et al., 2023) integrate user instructions into UI variants, enhancing nuanced user-LLM interaction. Gao et al. (Gao et al., 2024a) further identified key phases and interaction modes in Human-LLM interaction.
In the visual tools domain, the integration of LLMs with models like HuggingGPT(Shen et al., 2023) and visual foundation models(Ramesh et al., 2021; AI, 2022, 2023) has marked significant progress. However, there remains a gap in fully leveraging the combined potential of visual models and LLMs for improving design generation, particularly in interactive UI design. Our work addresses this by introducing a divide-and-conquer approach to natural language descriptions as a design-specific Domain-Specific Language (DSL) that is easily interpretable by both humans and LLMs. This DSL allows for detailed descriptions of global and local design elements, enhancing prompt clarity.
This form of prompt augmentation functions as a “software tool,” enabling designers to interact more efficiently with the system. By focusing less on crafting complex prompts and more on the core content of the design prototype, designers can actively participate in the design process. This provides real-time control over intermediate steps, allowing them to adjust the design direction, thereby improving the quality and coherence of the final output.
@@ 3. Formative Study and Findings
In this section, we conducted a study to gain insights from a professional designer’s perspective on the current design process when working with AI-powered design tools, identify the challenges UI designers face and explore potential improvements in design tools to better support their workflows.

### v3

# Towards Human-AI Synergy in UI Design: Leveraging LLMs for UI Generation with Intent Clarification and Alignment

@@ Abstract.
In automated UI design generation, a key challenge is the lack of support for iterative processes, as most systems focus solely on end-to-end output. This stems from limited capabilities in interpreting design intent and a lack of transparency for refining intermediate results. To better understand these challenges, we conducted a formative study that identified concrete and actionable requirements for supporting iterative design. Guided by these findings, we propose PrototypeFlow, a human-centered, multimodal-driven system for automated UI generation. PrototypeFlow takes natural language descriptions and layout preferences as input to generate the high-fidelity UI design. At its core is a theme design module that clarifies implicit design intent through prompt enhancement and orchestrates sub-modules for component-level generation. Designers retain full control over inputs, intermediate results, and final prototypes—enabling flexible, targeted refinement by steering generation and directly editing outputs. Our experiments and user studies confirmed the effectiveness and usefulness of our proposed PrototypeFlow.
Interactive design, User interface, Large language models, Image generation, Design Intent Clarification, Intent-Design Alignment, Automated Prototype Generation
copyright: nonecopyright: acmcopyrightjournalyear: 2022doi: XXXXXXX.XXXXXXXbooktitle: The 37th Annual ACM Symposium on User Interface Software and Technology(UIST’ 24)isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and tools
Figure 1. Overview of the user interface (UI) prototype creation process facilitated by PrototypeFlow, our multi-module collaborative generation system. (A) Designers input design intentions through prompts and initial UI layouts. (B) Our multi-modal generation system produces detailed prototypes. (C) These prototypes support top-down refinement from themes to specific components, with the option to regenerate at each level. (D) The output is a high-fidelity prototype, both visually rich and functionally grounded.
@@ 1. Introduction
To gain deeper insight into human-AI collaboration challenges in AI-driven design tools, we interviewed 10 professional UI/UX designers. This study revealed five key shortcomings: (F1) Need for streamlined design workflows supported by trend- and brand-aligned knowledge; (F2) Need for more input control and flexible output editability in the design generation process; (F3) Need for better support for expressing design intent through prompts; (F4) Need for precise control in generation processes; and (F5) Need for maintaining thematic consistency and coherence across generated components.
@@
In response to these observations, we introduce PrototypeFlow, a novel interactive design system that enables multi-module collaboration for UI prototype generation. This system empowers designers to craft high-fidelity prototypes that are editable, customisable and comprehensible at each stage of the generation process. As depicted in Fig. 1, PrototypeFlow begins with two design inputs: a textual description and a wireframe layout (Fig. 1 (A)). The text outlines the general design requirements (e.g., “Starting page for an intelligent design assistant”), while the wireframe offers a preliminary UI layout, enabling designers to convey their initial and nuanced design concepts.
@@
In the prototype generation phase (Fig. 1(B)), PrototypeFlow employs a top-down strategy, beginning with a central module that establishes the overarching theme and performs design intent enhancement through prompt augmentation. Following the decoupling approach, the central module coordinates the automatic, iterative creation of components, leveraging a cache pool to seamlessly integrate details from various sub-modules, each specially designed with expertise in different aspects. These specialized modules, including a text content generation module, image content generation module and retrieval-based icon module, ensure alignment with the overall design. This integrated approach allows PrototypeFlow to maintain aesthetic consistency, elevating the quality and coherence of the final prototype.
@@
In the review and editing phase (Fig. 1 (C)), PrototypeFlow offers deeper insights and detailed control within its interactive environment. When the user is dissatisfied with the theme color and the generated image, she directly changes the theme from green to yellow and revises the image description from “Flat style illustration with green…” to “a woman pointing at a robot to…”. Because the theme prompt is a global parameter, this triggers regeneration of the entire prototype. In comparison, for element-specific edits, the system supports localized updates. For instance, as shown in Figure 7(f), if the user only updates the content field of the second row (representing the image element) in the content editor and clicks “Regenerate Prototype”, only the image module is invoked to regenerate that particular content. It facilitates precise, iterative refinement, balancing automated generation with human customisation for efficient and personalized integration. Additionally, the generated prototypes (Fig. 1 (D)) can be saved in SVG or JSON formats, highlighting the system’s practicality and supporting high-quality creative output.
@@
For evaluation, we first conducted two automated quantitative evaluations, which validate PrototypeFlow’s ability to generate realistic and detailed UI designs. Our ablation study further highlights the significant impact and necessity of each module. In addition to the automated evaluations, we carried out three user studies involving 16 participants to assess the perceived usefulness of our work. The findings suggest that PrototypeFlow was positively received and showed significant potential in addressing the five identified challenges. Our work includes three main contributions:
@@
  • To the best of our knowledge, this work is among the first to report findings from semi-structured interviews with professional UI/UX designers working with GenAI design tools, detailing their current workflows and identifying five key shortcomings.
  • We introduce PrototypeFlow, a modular and interactive UI design system that enables efficient, user-centered prototype generation. The system coordinates specialized modules to regenerate the entire prototype in response to high-level changes, such as edits to the layout or theme, and also allows rapid, precise updates for component-level regeneration.
  • We conducted extensive experiments and user studies, which confirms the effectiveness of PrototypeFlow and surfaces actionable insights into how GenAI systems can better align with designers’ mental models and communication styles to support their creative workflow.
@@ 2. Related Work
Adobe Illustrator Figma / Adobe XD Midjourney / Stable Diffusion Vercel’s V0 Uizard PrototypeFlow (Our)
@@
Prompt Enhancement N/A N/A Not Supported Not Supported Not Supported Supported
Editable Checkpoints in Generation N/A N/A Not Supported Not Supported Not Supported Supported
Editable Theme Generation with Downstream Control N/A N/A Not Supported Not Supported Not Supported Supported
@@ 2.1. GUI Design Tools and Techniques
For creative design inspiration, text-based GUI retrieval, such as Guigle(Bernal-Cárdenas et al., 2019), leverages automated crawling and natural language processing to perform efficient searches through app hierarchies. Systems like Gallery D.C.(Chen et al., 2019), GUI2WiRe(Kolthoff et al., 2020), and RaWi(Kolthoff et al., 2023a) have further enhanced component extraction from screenshots, enabling flexible search based on dimensions, color, and text. Visual and multi-modal retrieval methods extend this by incorporating richer inputs such as screenshots and wireframes. Methods like WAE(Chen et al., 2020a), VINS(Bunian et al., 2021), and Swire(Huang et al., 2019) bridge the gap between low and high-fidelity designs, while methods such as WireGen(Feng et al., 2023) focus on automatically generating wireframes to connect low- and mid-fidelity stages, and Screen2Vec(Li et al., 2021) introduces multi-modal embeddings for diverse GUI content. In high-fidelity prototyping, tools like GUIGAN(Zhao et al., 2021) and research by Forrest et al.(Huang et al., 2021) focus on component retrieval and arrangement for detailed prototypes. However, these retrieval-based methods still rely heavily on existing designs, limiting creative flexibility. They address early-stage inspiration but leave much of the manual effort required for design refinement and customization.
@@
On the generative front, techniques like Layout2Image(Biswas et al., 2021), VAE(Bengio and LeCun, 2014), MidJourney(Midjourney, 2024), and Stable Diffusion(Rombach et al., 2022b) offer new possibilities for design generation. However, these methods often result in unstructured, non-editable outputs, making them challenging for GUI design. Projects such as PLay(Cheng et al., 2023) and DocSynth(Biswas et al., 2021) have made contributions to layout generation, facilitating the creation of low-fidelity prototypes. However, they provide limited support for detailed component creation or high-fidelity refinement. ndustry tools like Uizard(Uizard, 2024), Vercel’s V0(Vercel, 2024), and Figma plugins(uxp, 2024) provide automated prototype generation. Many of these tools, including Vercel and recent research systems, generate code as output (Beltramelli, 2018; Wu et al., 2024). However, these approaches primarily offer initial starting points, and the missing puzzle piece is how AI can continually collaborate with designers throughout the design iteration workflow, from concept to refinement.
While recent AI-powered design tools have made significant progress, they still present notable limitations in both input control and output editability. Most existing systems allow designers to provide either natural language prompts or image-based inputs, but rarely support multimodal input that combines wireframes with textual prompts. As a result, designers have limited ability to specify layout at the pixel level or to communicate detailed functional intent. In response to these gaps, our system introduces multimodal input, enabling designers to combine wireframes and natural language prompts for precise layout specification, functional accuracy, and explicit communication of design intent. For output, many previous approaches generate static images or code outputs (Beltramelli, 2018; Wu et al., 2024), which can be difficult to refine, especially for those without programming experience. By contrast, our system produces editable SVG prototypes. This allows designers to directly manipulate visual properties and efficiently iterate, bridging the gap between generative automation and practical, high-fidelity design work. A detailed comparison of input and output modalities across leading tools is presented in Table 1.
2.2. Human-AI Interaction in GUI
Large Language Models (LLMs) have significantly advanced human-computer interaction, especially in graphical user interface (GUI) contexts, by enabling natural language-driven workflows and automating design processes. For instance, Wang et al.(Wang et al., 2023a), Widget Captioning(Li et al., 2020b) and Stylette(Kim et al., 2022) showcase how LLMs facilitate intuitive conversational and command-based interactions in mobile and web UIs. Other works, such as MenuCraft (Kargaran et al., 2023), SUGILITE (Li et al., 2017), and Duan et al.(Duan et al., 2024), further demonstrate the versatility of LLMs in automating menu generation, task execution, and providing automated feedback on UI designs. Collaborative and educational applications are also emerging, with systems like CollabCoder(Gao et al., 2024b) and VIVID (Choi et al., 2024) highlighting LLMs’ roles in team analysis and generating dialogues from educational content.
@@ 2.2. Human-AI Interaction in GUI
While prior work has discussed broad HCI challenges in UI/UX, such as ethical considerations and fragmented tool ecosystems (Li et al., 2024), as well as design workflows from ideation to mock-up (Lu et al., 2022). There remains limited research on how GenAI can robustly support iterative and collaborative GUI design in professional practice. Some recent studies have begun to explore modules that clarify intents during the process of human-AI collaboration through interactions between end users and AI using natural language instructions (Li et al., 2018; Srivastava et al., 2017). SOVITE (Li et al., 2020a) expanded on this by enhancing system transparency through a mutual disambiguation pattern (Oviatt, 1999), where inputs from one modality help clarify inputs from another, allowing for breakdown repair. However, interactive design generation remains under-explored and often lacks reliability and clarity. Motivated by these gaps, our work introduces a decoupled generation approach with transparent, editable checkpoints throughout the generation process, allowing both designers and AI to collaboratively refine and clarify intermediate results (see Table 1).
2.3. Large Language Models for GUI Generation
Recent advances in large language models (LLMs) have opened new directions for automated GUI generation. Techniques such as chain-of-thought prompting and curriculum-driven task automation (Yao et al., 2022; Gravitas, 2023; Lu et al., 2024; Wang et al., 2023c) allow LLMs to complete tasks with minimal human intervention. Integrating LLMs with visual models (e.g., HuggingGPT (Shen et al., 2023), Stable Diffusion (AI, 2022, 2023)) further enables multi-modal workflows that bridge text and images.
While general prompt enhancement methods (Suzgun and Kalai, 2024; Wang et al., 2023b) have improved LLM performance across a broad range of tasks, prior work on GUI generation often focuses on static prompt-based generation or one-off design variants (Dang et al., 2023; Gao et al., 2024a), with limited attention to maintaining design coherence, enabling iterative refinement, or supporting transparent design workflows.
Our work grounds these general advances in the context of GUI design generation by introducing a divide-and-conquer approach that translates natural language into a domain-specific language (DSL) for editable component-level specification. We then support automatic prompt enhancement and editable theme generation, enabling theme-level control and allowing designers to refine outputs for consistency in color, style, usage and layout. Finally, we employ an LLM-based controller to ensure transparency and real-time edits, thus overcoming the static and non-transparent workflows of previous approaches (see Section 3, Table 1).
@@ 3. Formative Study and Findings
In this section, we conducted a study to gain insights from a professional designer’s perspective on the current design process when working with AI-powered design tools, identify the challenges UI designers face and explore potential improvements in design tools to better support their workflows. While prior studies—such as Li et al. (Li et al., 2024), which analyzes a broad range of issues from ethical concerns to fragmented tool ecosystems, and Lu et al. (Lu et al., 2022), which examines workflows related to inspiration search, mock-up generation, and iterative styling—have contributed important insights about the overall design process, there remains a gap in understanding how GenAI tools can directly and practically support iterative design refinement in real-world professional contexts. Most existing research highlights the importance of iteration and encourages for improved support, but often addresses broad, end-to-end workflows or “create from scratch” scenarios without actionable and fine-grained insights. By contrast, our study focuses specifically on the concrete, collaborative phase of design: when objectives are already well-defined, teams must follow to established design systems or branding, and the designer’s role is to turn explicit requirements into production-ready prototypes.
By grounding our interviews in these real-world, team-based contexts, our findings uncover some concrete challenges and needs that are not addressed in earlier, more general studies. We contribute actionable guidance on how Generative tools can better balance accuracy, creative flexibility, and seamless integration with existing workflows.